OpenStudy (anonymous):

How many units apart are the points (5, 20) and (10, 8)?

OpenStudy (amistre64):

if we move the points so that one of them is at the origin; we should be able to tell how far over and up the other point is and can use the pythag thrm to obtain a result

OpenStudy (lgbasallote):

i prefer distance formula but some people dont like memorizing so learning that is a good idea too

OpenStudy (amistre64):

yeah, and that is all the distance formula does; it moves the points and pythags it :)

OpenStudy (amistre64):

(5, 20) and (10, 8) -5-20 -5 -20 ----------------- 0 , 0 5 , -12 d^2 = 5^2 + 12^2 d = sqrt(25 + 144)

OpenStudy (anonymous):

169?

OpenStudy (amistre64):

sqrt(169) but yes

OpenStudy (anonymous):

13. (: thanks so much!

OpenStudy (amistre64):

yw